John is twice as old as Tom. In three years Tom will be 11 years old. How old will John be in three years?
Answer by ankor@dixie-net.com(22740)

Let J = John's age now
Let T = Tom's age now
:
Write an equation for each statement:
:
"John is twice as old as Tom"
J = 2T
:
" In three years Tom will be 11 years old."
T + 3 = 11
T = 11 - 3
T = 8 yrs old now
:
How old will John be in three years?
:
Find J's age now:
J = 2T, therefore:
J = 2(8)
J = 16 yrs old now
:
In 3 yrs, John would be 19 yrs old, obviously
;
:
That's pretty simple, right?